OpenStudy (anonymous):

Quadrilateral LMNO is a rectangle. Find MN. A. 7 B. 10 C. 18 D. 27

OpenStudy (anonymous):

Is there a picture?

OpenStudy (anonymous):

i got x=9 but it doesnt seem to fit

OpenStudy (anonymous):

4x-1=5x-8 -1=x-8 7=x

OpenStudy (anonymous):

therefore MN = (5 x 7) - 8 = 27